On a iq test with a mean of 100 and a standard deviation of 16, at approximately what percentile will an iq if 104 falls
Answer #1

Mu=100, sigma=16, x=104

P(x<104) = P( Z < (104 - mu )/sigma ) = P(Z < (104-100)/16 )

= P(Z < 0.25 ) = 0.5987

Approximately 60th percentile will an iq if 104 calls.